### v1

# The Mechanistic Emergence of Symbol Grounding in Language Models

Abstract
Symbol grounding (Harnad, 1990) describes how symbols such as words acquire their meanings by connecting to real-world sensorimotor experiences. Recent work has shown preliminary evidence that grounding may emerge in (vision-)language models trained at scale without using explicit grounding objectives. Yet, the specific loci of this emergence and the mechanisms that drive it remain largely unexplored. To address this problem, we introduce a controlled evaluation framework that systematically traces how symbol grounding arises within the internal computations through mechanistic and causal analysis. Our findings show that grounding concentrates in middle-layer computations and is implemented through the aggregate mechanism, where attention heads aggregate the environmental ground to support the prediction of linguistic forms. This phenomenon replicates in multimodal dialogue and across architectures (Transformers and state-space models), but not in unidirectional LSTMs. Our results provide behavioral and mechanistic evidence that symbol grounding can emerge in language models, with practical implications for predicting and potentially controlling the reliability of generation.
1 Introduction
Symbol grounding (Harnad, 1990) refers to the problem of how abstract and discrete symbols, such as words, acquire meaning by connecting to perceptual or sensorimotor experiences. Extending to the context of multimodal machine learning, grounding has been leveraged as an explicit pre-training objective for vision-language models (VLMs), by explicitly connecting linguistic units to the world that gives language meanings (Li et al., 2022; Ma et al., 2023). Through supervised fine-tuning with grounding signals, such as entity-phrase mappings, modern VLMs have achieved fine-grained understanding at both region (You et al., 2024; Peng et al., 2024; Wang et al., 2024) and pixel (Zhang et al., 2024b; Rasheed et al., 2024; Zhang et al., 2024a) levels.
With the rising of powerful autoregressive language models (LMs; OpenAI, 2024; Anthropic, 2024; Comanici et al., 2025, inter alia) and their VLM extensions, there is growing interest in identifying and interpreting their emergent capabilities. Recent work has shown preliminary correlational evidence that grounding may emerge in LLMs (Sabet et al., 2020; Shi et al., 2021; Wu et al., 2025b) and VLMs (Cao et al., 2025; Bousselham et al., 2024; Schnaus et al., 2025) trained at scale, even when solely optimized with the simple next-token prediction objective. However, the potential underlying mechanisms that lead to such an emergence are not well understood. To address this limitation, our work seeks to understand the emergence of symbol grounding in LMs, causally and mechanistically tracing how symbol grounding arises within the internal computations.
We begin by constructing a minimal testbed, motivated by the annotations provided in the CHILDES corpora (MacWhinney, 2000), where child–caregiver interactions provide cognitively plausible contexts for studying symbol grounding alongside verbal utterances. In our framework, each word is represented in two distinct forms: one token that appears in non-verbal scene descriptions (e.g., a box in the environment) and another that appears in spoken utterances (e.g., box in dialogue). We refer to these as environmental tokens (ENV) and linguistic tokens (LAN), respectively. A deliberately simple word-level tokenizer assigns separate vocabulary entries to each form, ensuring that they are treated as entirely different tokens by the language model. This framework enforces a structural separation between scenes and symbols, preventing correspondences from being reduced to trivial token identity. Under this setup, we can evaluate whether a model trained from scratch is able to predict the linguistic form from its environmental counterpart.
We quantify the level of grounding using surprisal: specifically, we compare how easily the model predicts a linguistic token (LAN) when its matching environmental token (ENV) is present versus when unrelated cues are given instead. A lower surprisal in the former condition indicates that the model has learned to align environmental grounds with linguistic forms. We find that LMs do learn to ground: the presence of environmental tokens consistently reduces surprisal for their linguistic counterparts, in a way that simple co-occurrence statistics cannot fully explain. To study the underlying mechanisms, we apply saliency analysis (Wang et al., 2023) and the tuned lens (Belrose et al., 2023), which converge on the result that grounding relations are concentrated in the middle layers of the network. Further analysis of attention heads reveals patterns consistent with the aggregate mechanism (Bick et al., 2025), where attention heads support the prediction of linguistic forms by retrieving their environmental grounds in the context.
Finally, we demonstrate that these findings generalize beyond the minimal CHILDES data and Transformer models. They appear in a multimodal setting with the Visual Dialog dataset (Das et al., 2017), and in state-space models (SSMs) such as Mamba-2 (Dao & Gu, 2024). In contrast, we do not observe grounding in unidirectional LSTMs, consistently with their sequential state compression and lack of content-addressable retrieval. Taken together, our results show that symbol grounding can mechanistically emerge in autoregressive LMs, while also delineating the architectural conditions under which it can arise.
2 Related Work
2.1 Language grounding.
Referential grounding has long been framed as the lexicon acquisition problem: how words map to referents in the world (Harnad, 1990; Gleitman & Landau, 1994; Clark, 1995). Early work focused on word-to-symbol mappings, designing learning mechanisms that simulate children’s lexical acquisition and explain psycholinguistic phenomena (Siskind, 1996; Regier, 2005; Goodman et al., 2007; Fazly et al., 2010). Subsequent studies incorporated visual grounding, first by aligning words with object categories (Roy & Pentland, 2002; Yu, 2005; Xu & Tenenbaum, 2007; Yu & Ballard, 2007; Yu & Siskind, 2013), and later by mapping words to richer visual features (Qu & Chai, 2010; Mao et al., 2019; 2021; Pratt et al., 2020). More recently, large-scale VLMs trained with paired text–image supervision have advanced grounding to finer levels of granularity, achieving region-level (Li et al., 2022; Ma et al., 2023; Chen et al., 2023; You et al., 2024; Wang et al., 2024) and pixel-level (Xia et al., 2024; Rasheed et al., 2024; Zhang et al., 2024b) grounding, with strong performance on referring expression comprehension (Chen et al., 2024a).
Recent work suggests that grounding emerges as a property of VLMs trained without explicit supervision, with evidence drawn from attention-based spatial localization (Cao et al., 2025; Bousselham et al., 2024) and cross-modal geometric correspondences (Schnaus et al., 2025). However, all prior work focused exclusively on static final-stage models, overlooking the training trajectory, a crucial aspect for understanding when and how grounding emerges. In addition, existing work has framed grounding through correlations between visual and textual signals, diverging from the definition by Harnad (1990), which emphasizes causal links from symbols to meanings. To address these issues, we systematically examine learning dynamics throughout the training process, applying causal interventions to probe model internals and introducing control groups to enable rigorous comparison.
2.2 Emergent capabilities and learning dynamics of LMs.
A central debate concerns whether larger language models exhibit genuinely new behaviors: Wei et al. (2022) highlight abrupt improvements in tasks, whereas later studies argue such effects are artifacts of thresholds or in-context learning dynamics (Schaeffer et al., 2023; Lu et al., 2024). Beyond end performance, developmental analyses show that models acquire linguistic abilities in systematic though heterogeneous orders with variability across runs and checkpoints (Sellam et al., 2021; Blevins et al., 2022; Biderman et al., 2023; Xia et al., 2023; van der Wal et al., 2025). Psychology-inspired perspectives further emphasize controlled experimentation to assess these behaviors (Hagendorff, 2023), and comparative studies reveal both parallels and divergences between machine and human language learning (Chang & Bergen, 2022; Evanson et al., 2023; Chang et al., 2024; Ma et al., 2025). At a finer granularity, hidden-loss analyses identify phase-like transitions (Kangaslahti et al., 2025), while distributional studies attribute emergence to stochastic differences across training seeds (Zhao et al., 2024). Together, emergent abilities are not sharp discontinuities but probabilistic outcomes of developmental learning dynamics. Following this line of work, we present a probability- and model internals–based analysis of how symbol grounding emerges during language model training.
2.3 Mechanistic interpretability of LMs.
Mechanistic interpretability has largely focused on attention heads in Transformers (Elhage et al., 2021; Olsson et al., 2022; Meng et al., 2022; Bietti et al., 2023; Lieberum et al., 2023; Wu et al., 2025a). A central line of work established that induction heads emerge to support in-context learning (ICL; Elhage et al., 2021; Olsson et al., 2022), with follow-up studies tracing their training dynamics (Bietti et al., 2023) and mapping factual recall circuits (Meng et al., 2022). At larger scales, Lieberum et al. (2023) identified specialized content-gatherer and correct-letter heads, and Wu et al. (2025a) showed that a sparse set of retrieval heads is critical for reasoning and long-context performance. Relatedly, Wang et al. (2023) demonstrated that label words in demonstrations act as anchors: early layers gather semantic information into these tokens, which later guide prediction. Based on these insights, Bick et al. (2025) proposed that retrieval is implemented through a coordinated gather-and-aggregate (G&A) mechanism: some heads collect content from relevant tokens, while others aggregate it at the prediction position. Other studies extended this line of work by analyzing failure modes and training dynamics (Wiegreffe et al., 2025) and contrasting retrieval mechanisms in Transformers and SSMs (Arora et al., 2025). Whereas prior analyses typically investigate ICL with repeated syntactic or symbolic formats, our setup requires referential alignment between linguistic forms and their environmental contexts, providing a complementary testbed for naturalistic language grounding.
3 Method
| Dataset | Training Example | Test Example | |||
[…]
| Image-Grounded Dialogue |
[…]
tticblue!10
Q can you tell what book it’s reading A the marriage of true minds by stephen evans |
[…]
tticblue!10
[…]
tticblue!10
what do we have here? |
3.1 Dataset and Tokenization
To capture the emergent grounding from multimodal interactions, we design a minimal testbed with a custom word-level tokenizer, in which every lexical item is represented in two corresponding forms: one token that appears in non-verbal descriptions (e.g., a book in the scene description) and another that appears in utterances (e.g., book in speech). We refer to these by environmental (ENV) and linguistic tokens (LAN), respectively. For instance, book and book are treated as distinct tokens with separate integer indices; that is, the tokenization provides no explicit signal that these tokens are related, so any correspondence between them must be learned during training rather than inherited from their surface form. We instantiate this framework in three datasets, ranging from child-directed speech transcripts to image-based dialogue.
Child-directed speech. The Child Language Data Exchange System (CHILDES; MacWhinney, 2000) provides transcripts enriched with environmental annotations.‡‡‡See the manual for data usage: https://talkbank.org/0info/manuals/CHAT.pdf We use the spoken utterances as the linguistic tokens (LAN) and the environmental descriptions as the environment tokens (ENV). The environmental context is drawn from three annotation types:
Local events: simple events, pauses, long events, or remarks interleaved with the transcripts.
Action tiers: actions performed by the speaker or listener (e.g., %act: runs to toy box). These also include cases where an action replaces speech (e.g., 0 [% kicks the ball]).
Situational tiers: situational information tied to utterances or to larger contexts (e.g., %sit: dog is barking).
Caption-grounded dialogue. The Visual Dialog dataset (Das et al., 2017) pairs MSCOCO images (Lin et al., 2014) with sequential question-answering based multi-turn dialogues that exchange information about each image. Our setup uses MSCOCO captions as the environmental tokens (ENV) and the dialogue turns form the linguistic tokens (LAN). In this pseudo cross-modal setting, textual descriptions of visual scenes ground natural conversational interaction. Compared to CHILDES, this setup introduces richer semantics and longer utterances, while still using text-based inputs for both token types, thereby offering a stepping stone toward grounding in fully visual contexts.
Image-grounded dialogue. To move beyond textual proxies, we consider an image-grounded dialogue setup, using the same dataset as the caption-grounded dialogue setting. Here, a frozen vision transformer (ViT; Dosovitskiy et al., 2020) directly tokenizes each RGB image into patch embeddings, with each embedding treated as an ENV token, analogously to the visual tokens in modern VLMs. We use DINOv2 (Oquab et al., 2024) as our ViT tokenizer, as it is trained purely on vision data without auxiliary text supervision (in contrast to models like CLIP; Radford et al., 2021), thereby ensuring that environmental tokens capture only visual information. The linguistic tokens (LAN) remain unchanged from the caption-grounded dialogue setting, resulting in a realistic multimodal interaction where conversational utterances are grounded directly in visual input.
3.2 Evaluation Protocol
We assess symbol grounding with a contrastive test that asks whether a model assigns a higher probability to the correct linguistic token when the matching environmental token is in context, following the idea of priming in psychology. This evaluation applies uniformly across datasets (Table 1): in CHILDES and caption-grounded dialogue, environmental priming comes from descriptive contexts; in image-grounded dialogue, from ViT-derived visual tokens. We compare the following conditions:
Match (experimental condition): The context contains the corresponding ENV token for the target word, and the model is expected to predict its LAN counterpart.
Mismatch (control condition): The context is replaced with a different ENV token. The model remains tasked with predicting the same LAN token; however, in the absence of corresponding environmental cues, its performance is expected to be no better than chance.
For example (first row in Table 1), when evaluating the word , the input context is
where the model is expected to predict for the blank, and the role token CHI indicates the involved speaker or actor’s role being a child. In the control (mismatch) condition, the environmental token box is replaced by another valid noun such as toy.
Context templates. For a target word with linguistic token and environmental token , we denote as a set of context templates of . For example, when , a can be
where [FILLER] is to be replaced with an environmental token, and the blank indicates the expected prediction as in Eq. (1). In the match condition, the context is constructed by replacing [FILLER] with in . In the mismatch condition, the context uses as the filler, while the prediction target remains .
For the choices of and , we construct the vocabulary with 100 nouns from the MacArthur–Bates Communicative Development Inventories (Fenson et al., 2006) that occur frequently in our corpus. Each word serves once as the target, with the remaining used to construct mismatched conditions. For each word, we create context templates, which contain both ENV and LAN tokens. Details of the vocabulary and context template construction can be found in the Appendix A.
Grounding information gain. Following prior work, we evaluate how well an LM learns a word using the mean surprisal over instances. The surprisal of a word given a context is defined as where denotes the probability, under an LM parameterized by , that the next word is conditioned on the context . Here, quantifies the unexpectedness of predicting , or the pointwise information carried by conditioned on the context.
The grounding information gain for is defined as
This is a sample-based estimation of the expected log-likelihood ratio between the match and mismatch conditions
which quantifies how much more information the matched ground provides for predicting the linguistic form, compared to a mismatched one. A positive indicates that the matched environmental token increases the predictability of its linguistic form. We report , and track across training steps to analyze how grounding emerges over time.
3.3 Model Training
We train LMs from random initialization, ensuring that no prior linguistic knowledge influences the results. Our training uses the standard causal language modeling objective, as in most generative LMs. To account for variability, we repeat all experiments with 5 random seeds, randomizing both model initialization and corpus shuffle order. Our primary architecture is Transformer (Vaswani et al., 2017) in the style of GPT-2 (Radford et al., 2019) with 18, 12, and 4 layers, with all of them having residual connections. We extend the experiments to 4-layer unidirectional LSTMs (Hochreiter & Schmidhuber, 1997) with no residual connections, as well as 12- and 4-layer state-space models (specifically, Mamba-2; Dao & Gu, 2024). For fair comparison with LSTMs, the 4-layer Mamba-2 models do not involve residual connections, whereas the 12-layer ones do. For multimodal settings, while standard LLaVA (Liu et al., 2023) uses a two-layer perceptron to project ViT embeddings into the language model, we bypass this projection in our case and directly feed the DINOv2 representations into the LM. We obtain the developmental trajectory of the model by saving checkpoints at various training steps, sampling more heavily from earlier steps, following Chang & Bergen (2022).
4 Behavioral Evidence
4.1 Behavioral Evidence of Emergent Grounding
In this section, we ask: Does symbol grounding emerge behaviorally in autoregressive LMs? We first test whether models show systematic surprisal reduction when predicting a linguistic token when its environmental counterpart is in context (Figure 2, where the gap between the lines represent the grounding information gain). For Transformers (Figures 2(a) and 2(b)) and Mamba-2 (Figure 2(c)), surprisal in the match condition decreases steadily while that in the mismatch condition enters a high-surprisal plateau early, indicating that the models leverage environmental context to predict the linguistic form. In contrast, the unidirectional LSTM (Figure 2(d)) shows little separation between the conditions, reflecting the absence of grounding. Overall, these results provide behavioral evidence of emergent grounding: in sufficiently expressive architectures (Transformers and Mamba-2), the correct environmental context reliably lowers surprisal for its linguistic counterpart, whereas LSTMs fail to exhibit this effect, marking an architectural boundary on where grounding can emerge.
4.2 Behavioral Effects Beyond Co-occurrence
A natural concern is that the surprisal reductions might be fully explainable by shallow statistics: the models might have simply memorized frequent co-occurrences of ENV and LAN tokens, without learning a deeper and more general mapping. We test this hypothesis by comparing the tokens’ co-occurrence with the grounding information gain in the child-directed speech data.
We define co-occurrence between the corresponding ENV and LAN tokens at the granularity of a 512-token training chunk. For each target word , we count the number of chunks in which both its ENV and LAN tokens appear. Following standard corpus-analysis practice, these raw counts are log-transformed. For each model checkpoint, we run linear regression between the log co-occurrence and the grounding information gain of words, obtaining an statistic as a function of training time.
Figure 3 shows the values (orange) alongside the grounding information gain (blue) for different architectures. In both the Transformer and Mamba-2, rises sharply at the early steps but then goes down, even if the grounding information gain continues increasing. These results suggest that grounding in Transformers and Mamba-2 cannot be fully accounted for by co-occurrence statistics: while models initially exploit surface co-occurrence regularities, later improvements in grounding diverge from these statistics, indicating reliance on richer and more complicated features acquired during training. In contrast, LSTM shows persistently increasing but little increase in grounding information gain over training steps, suggesting that it encodes co-occurrence but lacks the architectural mechanism to transform it into predictive grounding.
4.3 Visual Dialogue with Captions and Images
We next test whether the grounding effects observed in CHILDES generalize to multimodal dialogue, using the Visual Dialog dataset. In this setting, the environmental ground is supplied either by captions or by image features (Table 1). For caption-grounded dialogue, the mismatch context is constructed in the same way as for CHILDES (Equation 2). For image-grounded dialogue, mismatch contexts are generated via Stable Diffusion 2 (Rombach et al., 2022)–based image inpainting, which re-generates the region defined by the ground-truth mask corresponding to the target word’s referent.
We train 12-layer Transformers with 5 random seeds. Similarly as Figures 2(a)–2(b) and Figures 3(a)–3(b), when captions serve as the environmental ground, Transformers show a clear surprisal gap between match and mismatch conditions (Figure 4(a)), with the grounding information gain increasing steadily while peaks early and declines (Figure 4(c)). Directly using image as grounds yields the same qualitative pattern (Figures 4(b) and 4(d)), although the observed effect is smaller. Both settings confirm that emergent grounding cannot be fully explained by co-occurrence statistics.
Overall, our findings demonstrate that Transformers are able to exploit environmental grounds in various modalities to facilitate linguistic prediction. The smaller but consistent gains in the image-grounded case suggest that while grounding from visual tokens is harder, the same architectural dynamics identified in textual testbeds still apply.
5 Mechanistic Explanation
In this section, we provide a mechanistic and interpretable account of the previous observation. We focus on a 12-layer Transformer trained on CHILDES with 5 random seeds, and defer broader generalization to the discussion.
5.1 The Emergence of Symbol Grounding
To provide a mechanistic account of symbol grounding, i.e., when it emerges during training and how it is represented in the network, we apply two interpretability analyses.
Saliency flow. For each layer , we compute a saliency matrix following Wang et al. (2023): , where denotes the attention matrix of head in layer . Each entry of quantifies the contribution of the corresponding attention weight to the cross-entropy loss , averaged across heads. Our analysis focuses on ground-to-symbol connections, i.e., flows from environmental ground (ENV) tokens to the token immediately preceding (and predicting) their linguistic forms (LAN).
Probing with the Tuned Lens. We probe layer-wise representations using the Tuned Lens (Belrose et al., 2023), which trains affine projectors to map intermediate activations to the final prediction space while keeping the LM output head frozen.
Results. Ground-to-symbol saliency is weak in the early stages of training but rises sharply later, peaking in layers 7–9 (Figure 5(a)), suggesting that mid-layer attention plays a central role in establishing symbol–ground correspondences. In addition, Figure 5(b) shows that early layers remain poor predictors even at late training stages (e.g., after 20,000 steps), whereas surprisal begins to drop markedly from layer 7 at intermediate stages (step 10,000), suggesting a potential representational shift in the middle layers.
5.2 Hypothesis: Gather-and-Aggregate Heads Implement Symbol Grounding
Building on these results, we hypothesize that specific Transformer heads in the middle layers enable symbol grounding. To test this, we examine attention saliencies for selected heads (Figure 6). We find that several heads exhibit patterns consistent with the gather and aggregate mechanisms described by Bick et al. (2025): gather heads (e.g., Figures 6(a) and 6(b)) compress relevant information into a subset of positions, while aggregate heads (e.g., Figures 6(c) and 6(d)) redistribute this information to downstream tokens. In our setups, saliency often concentrates on environmental tokens such as train, where gather heads pool contextual information into compact, retrievable states. In turn, aggregate heads broadcast this information from environmental ground (trainENV) to the token immediately preceding the linguistic form, thereby supporting the prediction of train. Taking these observations together, we hypothesize that the gather-and-aggregate heads implement the symbol grounding mechanism.
| Ckpt. | Gather Head | Aggregate Head | Original | ||||||
| Avg. | Avg. | Interv. | Ctrl. | Avg. | Avg. | Interv. | Ctrl. | ||
| Count | Layer | Sps. | Sps. | Count | Layer | Sps. | Sps. | ||
| 500 | 0.00 | - | - | - | 0.07 | 8.74 | 9.34 | 9.34 | 9.34 |
| 5000 | 0.35 | 3.32 | 6.37 | 6.38 | 2.28 | 7.38 | 6.51 | 6.39 | 6.38 |
| 10000 | 3.26 | 3.67 | 5.25 | 5.32 | 5.09 | 7.28 | 5.86 | 5.29 | 5.30 |
| 20000 | 5.76 | 3.59 | 4.69 | 4.79 | 6.71 | 7.52 | 5.62 | 4.76 | 4.77 |
5.3 Causal Interventions of Attention Heads
We then conduct causal interventions of attention heads to validate our previous hypothesis.
Operational definition. We identify attention heads as gather or aggregate following these standards:
Gather head. An attention head is classified as a gather head if at least 30% of its total saliency is directed toward the environmental ground token from the previous ones.
Aggregate head: An attention head is classified as an aggregate head if at least 30% of its total saliency flows from the environmental ground token to the token immediately preceding the corresponding linguistic token.
Causal intervention methods. In each context, we apply causal interventions to the identified head types and their corresponding controls. Following Bick et al. (2025), interventions are implemented by zeroing out the outputs of heads. For the control, we mask an equal number of randomly selected heads in each layer, ensuring they do not overlap with the identified gather or aggregate heads.
| Thres. | Ckpt. | Aggregate Head | Original | |||
| Avg. | Avg. | Interv. | Ctrl. | |||
| Count | Layer | Sps. | Sps. | |||
| 70% | 20k | 32.30 | 7.78 | 9.96 | 9.95 | 9.21 |
| 100k | 35.63 | 7.71 | 9.42 | 8.84 | 8.24 | |
| 200k | 34.99 | 7.80 | 8.95 | 8.15 | 7.76 | |
| 300k | 34.15 | 7.76 | 8.96 | 8.11 | 7.69 | |
| 90% | 20k | 10.66 | 8.33 | 9.51 | 9.43 | 9.21 |
| 100k | 13.90 | 8.26 | 8.95 | 8.50 | 8.24 | |
| 200k | 13.47 | 8.46 | 8.41 | 7.88 | 7.76 | |
| 300k | 12.73 | 8.42 | 8.40 | 7.87 | 7.69 | |
Results and discussions. As training progresses, the number of both gather and aggregate heads increases (Table 2), suggesting that these mechanisms emerge over the course of learning. Causal interventions reveal a clear dissociation: zeroing out aggregate heads consistently produces significantly higher surprisal compared to controls, whereas the gather head interventions have no such effect. This asymmetry suggests that gather heads serve in a role less critical in our settings, where the input template is semantically light and the environmental evidence alone suffices to shape the linguistic form. Layer-wise patterns further support this division of labor: gather heads cluster in shallow layers (3-4), while aggregate heads concentrate in mid layers (7-8). This resonates with our earlier probing results, where surprisal reductions became prominent only from layers 7-9. Together, these findings highlight aggregate heads in the middle layers as the primary account of grounding in the model.
5.4 Generalization to Visual Dialogue with Images
We also conduct causal interventions of attention heads on the VLM model to further validate our previous hypothesis.
Operational definition. We identify attention heads as aggregate following this standard (We do not define gather head): An attention head is classified as an aggregate head if at least a certain threshold (70% or 90% in our experiment settings) of its total image patch to end saliency flows from the patches inside bounding box to the token immediately preceding the corresponding linguistic token.
Causal intervention methods. In each context, we apply causal interventions to the identified head types and their corresponding controls in the language backbone of the model. Similar to section 5.3, interventions are implemented by zeroing out a head’s outputs. For the control, we mask an equal number of randomly selected heads in each layer, ensuring they do not overlap with the identified gather or aggregate heads.
Results and discussions. As training progresses, the number of aggregate heads increases first and then becomes steady (Figure 7), suggesting that these mechanisms emerge over the course of learning. Causal interventions reveal that zeroing out aggregate heads consistently produces significantly higher surprisal rises compared to controls. The average layer also align with the saliency heatmap, also shown in Figure 7.
6 Discussions
Generalization to full-scale VLMs. As an additional case study, we extend our grounding-as-aggregation hypothesis to a full-scale VLM, LLaVA-1.5-7B (Liu et al., 2023). Even in this heavily engineered architecture, we identify many attention heads exhibiting aggregation behavior consistent with our earlier findings (Figure 1(b)), reinforcing the view that symbol grounding arises from specialized heads. At the same time, full-scale VLMs present additional complications. Models like LLaVA use multiple sets of visual tokens, including CLIP-derived embeddings that already encode language priors, and global information may be stored in redundant artifact tokens rather than object-centric regions (Darcet et al., 2024). Moreover, the large number of visual tokens (environmental tokens, in our setup) substantially increases both computational cost and the difficulty of isolating genuine aggregation heads. These factors make systematic identification and intervention at scale a nontrivial challenge. For these reasons, while our case study highlights promising evidence of grounding heads in modern VLMs, systematic detection and causal evaluation of such heads at scale remains an open challenge. Future work will need to develop computationally viable methods for (i) automatically detecting aggregation heads across diverse VLMs, and (ii) applying causal interventions to validate their role in grounding. Addressing these challenges will be crucial for moving from anecdotal case studies to a more principled understanding of grounding in modern VLMs.
The philosophical roots of grounding, revisited. Our findings highlight the need to sharpen the meaning of grounding in multimodal models. Prior work has often equated grounding with statistical correlations between visual and textual signals, such as attention overlaps or geometric alignments (Bousselham et al., 2024; Cao et al., 2025; Schnaus et al., 2025). While informative, such correlations diverge from the classic formulation by Harnad (1990), which requires symbols to be causally anchored to their referents in the environment. On the other extreme, Gubelmann (2024) argued that the symbol grounding problem does not apply to LLMs as they “are connectionist, statistical devices that have no intrinsic symbolic structure.” In contrast, we discover emergent symbolic structure as an intrinsic mechanistic property: one that can be traced along training, observed in the specialization of attention heads, and validated through causal interventions. This provides not only a practical diagnostic protocol that reveals when and how models genuinely tie symbols to meaning beyond surface-level correlations, but also challenges the view that grounding is philosophically irrelevant to systems without explicit symbolic structure.
Practical implications to LM hallucinations. Our findings have practical implications for improving the reliability of LM outputs: by identifying aggregation heads that mediate grounding between environmental and linguistic tokens, we provide a promising mechanism to detect model reliability before generation. Our findings echo a pathway to mitigate hallucinations by focusing on attention control: many hallucination errors stem from misallocated attention in intermediate layers (Jiang et al., 2025; Chen et al., 2024b). Such attention-level signals can serve as early indicators of overtrust or false grounding, motivating practical solutions like decoding-time strategies to mitigate and eventually prevent hallucination (Huang et al., 2024).
Acknowledgement
This work was supported in part by NSF IIS-1949634, NSF SES-2128623, NSERC RGPIN-2024-04395, the Weinberg Cognitive Science Fellowship to ZM, a Vector Scholarship to XL, and a Canada CIFAR AI Chair award to FS. The authors would like to thank Songlin Yang and Jing Ding for their valuable feedback.
Appendix A Dataset Details
A.1 Context Templates
We select the target tokens following the given procedure:
1.
Get a list of words, with their ENV and LAN frequency both greater than or equal to 100 in the CHILDES dataset;
2.
Get another list of nouns from CDI;
3.
Take intersection and select top 100 words (by frequency of their ENV token) as target token list.
In CHILDES, all contexts are created with gpt-4o-mini followed by human verification if the genrated contexts are semantically light. We adopt the following prompt:
In visual dialogue (caption version and VLM version), we pre-define 10 sets of templates for each version:
A.2 Word Lists
CHILDES and Visual Dialog (Text Only). [box, book, ball, hand, paper, table, toy, head, car, chair, room, picture, doll, cup, towel, door, mouth, camera, duck, face, truck, bottle, puzzle, bird, tape, finger, bucket, block, stick, elephant, hat, bed, arm, dog, kitchen, spoon, hair, blanket, horse, tray, train, cow, foot, couch, necklace, cookie, plate, telephone, window, brush, ear, pig, purse, hammer, cat, shoulder, garage, button, monkey, pencil, shoe, drawer, leg, bear, milk, egg, bowl, juice, ladder, basket, coffee, bus, food, apple, bench, sheep, airplane, comb, bread, eye, animal, knee, shirt, cracker, glass, light, game, cheese, sofa, giraffe, turtle, stove, clock, star, refrigerator, banana, napkin, bunny, farm, money]
Visual Dialog (VLM). [box, book, table, toy, car, chair, doll, door, camera, duck, truck, bottle, bird, elephant, hat, bed, dog, spoon, horse, train, couch, necklace, cookie, plate, telephone, window, pig, cat, monkey, drawer, bear, milk, egg, bowl, juice, ladder, bus, food, apple, sheep, bread, animal, shirt, cheese, giraffe, clock, refrigerator, accordion, aircraft, alpaca, ambulance, ant, antelope, backpack, bagel, balloon, barrel, bathtub, beard, bee, beer, beetle, bicycle, bidet, billboard, boat, bookcase, boot, boy, broccoli, building, bull, burrito, bust, butterfly, cabbage, cabinetry, cake, camel, canary, candle, candy, cannon, canoe, carrot, cart, castle, caterpillar, cattle, cello, cheetah, chicken, chopsticks, closet, clothing, coat, cocktail, coffeemaker, coin, cosmetics]
Appendix B Implementation Details
We outline the key implementation details in this section and provide links to the GitHub repositories:
Model Training: https://github.com/Mars-tin/TraBank
CHILDES Processing: https://github.com/Mars-tin/PyChildes
B.1 Checkpointing
We save 33 checkpoints in total for text-only experiments and 16 checkpoints for the VLM setting.
CHILDES and Visual Dialog (Text Only). We save the intermediate steps: [0, 150, 300, 500, 1000, 1500, 2000, 2500, 3000, 3500, 4000, 4500, 5000, 5500, 6000, 6500, 7000, 7500, 8000, 8500, 9000, 9500, 10000, 11000, 12000, 13000, 14000, 15000, 16000, 17000, 18000, 19000, 20000]
Visual Dialog (VLM). We save the intermediate steps: [10000, 20000, 40000, 60000, 80000, 100000, 120000, 140000, 160000, 180000, 200000, 220000, 240000, 260000, 280000, 300000]
B.2 Training details.
For the text-only Transformer, Mamba2, and LSTM models, we randomly initialize them from scratch. The training process is conducted five times, each with a different random seed (using seeds 42, 142, 242, 342, and 442, respectively). The batch size is 16.
For VLM models, we randomly initialize the language model backbone from scratch and keep the DINOv2 vision encoder frozen. The training process is conducted five times for 300k steps, each with a different random seed (using seed 42, 142, 242, 342, and 442, respectively).
All the models use a word-level tokenizer. A list of hyperparameters is shown below:
Transformer and LSTM Model.
model_max_length: 512
learning rate: 5e-5
learning rate schedule: linear
warmup_steps: 1000
hidden_size: 768
beta1: 0.9
beta2: 0.95
weight_decay: 0
batch_size: 16
grad_clip_norm: 1.0
Mamba2 Model.
model_max_length: 512
learning rate: 4e-4
learning rate schedule: linear
warmup_steps: 2000
hidden_size: 768
beta1: 0.9
beta2: 0.95
weight_decay: 0.4
batch_size: 16
grad_clip_norm: 1.0
VLM Model.
model_max_length: 1024
learning rate: 2e-5
learning rate schedule: cosine
warmup_steps: 9000
hidden_size: 768
beta1: 0.9
beta2: 0.95
weight_decay: 0
batch_size: 16
grad_clip_norm: 1.0
B.3 Computational resources.
Each Transformer, Mamba2, and LSTM model is trained on a single A40 GPU within 5 hours. For VLM models, training is conducted on 2 A40 GPUs over 15 hours, using a batch size of 8 per device.
Appendix C Addendum to Results
C.1 Behavioral Analysis
C.2 Mechanistic Analysis
After identifying the set of gather and aggregate heads for each context, we conduct an overtime analysis to determine the proportion of saliency to the total saliency, as illustrated in Figure 8.